Suppose you want to test the claim the the paired sample data given below come from a population for which the mean difference is μd=0. xy7866678183687180529154928286 Use a 0.01 significance level to find the following: (a) The mean value of the differnces d for the paired sample data d¯= (b) The standard deviation of the differences d for the paired sample data sd= (c) The t test statistic t= (d) The positive critical value t= (e) The negative critical value t= (f) Does the test statistic fall in the critical region? A. No B. Yes (g) Construct a 99% conficence interval for the population mean of all differences x−y. <μd
Answer #1

Computational Table:

X Y d=x-y d^2
78 66 12 144
67 81 -14 196
83 68 15 225
71 80 -9 81
52 91 -39 1521
54 92 -38 1444
82 86 -4 16
Total -77 3627

a)

b)

c) Test statistic:

d)

Degrees of freedom = n-1 = 7-1 = 6

Critical value:

...............................From t table

Therefore,

Positive Critical value: 3.7074

Negative Critical value: -3.7074

Conclusion:

Test statistic (t) lies in the rejection region,i.e. -3.7074 < -1.35 < 3.7074, That is Fail to Reject Ho at 1% level of significance.   

f) Yes, Test statistic -1.35 is Falls in the critical region

g) Construct 99% Confidence interval:

We are 99% Confidence that population mean difference is lies in that interval.